# Similarity-Dissimilarity Loss for Multi-label Supervised Contrastive Learning

### v4

@@ A.4 Proof of Theorem 4
Let us establish the strict inequality $\mathcal{K}_{4}^{s}\mathcal{K}_{4}^{d}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ through direct comparison. From definitions (14) and (15), we have:
$\mathcal{K}_{4}^{s}\mathcal{K}_{4}^{d}=\frac{|\tilde{\boldsymbol{y}}_{p_{4}}|}{|\tilde{\boldsymbol{y}}_{i}|}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}=\frac{|\tilde{\boldsymbol{y}}_{p_{3}}-\tilde{\boldsymbol{y}}_{p_{3}}^{d}|}{|\tilde{\boldsymbol{y}}_{i}|}\cdot\frac{1}{1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|}$
@@
$\frac{|\tilde{\boldsymbol{y}}_{p_{4}}|(1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|)}{|\tilde{\boldsymbol{y}}_{i}|(1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|)}>\frac{|\tilde{\boldsymbol{y}}_{p_{4}}-\tilde{\boldsymbol{y}}_{p_{3}}^{d}|}{|\tilde{\boldsymbol{y}}_{i}|(1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|)}$
@@
$|\tilde{\boldsymbol{y}}_{p_{4}}|(1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|)>|\tilde{\boldsymbol{y}}_{p_{3}}-\tilde{\boldsymbol{y}}_{p_{3}}^{d}|$
By the cardinality constraint (22) in the theorem:
$|\tilde{\boldsymbol{y}}_{p_{3}}|(1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|)>|\tilde{\boldsymbol{y}}_{p_{3}}-\tilde{\boldsymbol{y}}_{p_{3}}^{d}|$
@@ A.5 Proof of Theorem 5
From definitions (14) and (15), we have:
$\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ $=\frac{|\tilde{\boldsymbol{y}}_{p_{3}}-\tilde{\boldsymbol{y}}_{p_{3}}^{d}|}{|\tilde{\boldsymbol{y}}_{i}|}\cdot\frac{1}{1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|}$
$\mathcal{K}_{5}^{s}\mathcal{K}_{5}^{d}$ $=\frac{1}{1+|\tilde{\boldsymbol{y}}_{p_{5}}^{d}|}$
@@
$\frac{\mathcal{K}_{5}^{s}\mathcal{K}_{5}^{d}}{\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}}=\frac{|\tilde{\boldsymbol{y}}_{i}|(1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|)}{|\tilde{\boldsymbol{y}}_{p_{3}}-\tilde{\boldsymbol{y}}_{p_{3}}^{d}|(1+|\tilde{\boldsymbol{y}}_{p_{5}}^{d}|)}$
@@
$|\tilde{\boldsymbol{y}}_{p_{3}}-\tilde{\boldsymbol{y}}_{p_{3}}^{d}|\leq|\tilde{\boldsymbol{y}}_{i}|$
Given the constraint (24), $|\tilde{\boldsymbol{y}}_{p_{5}}^{d}|\leq|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|$, we have:
$\frac{|\tilde{\boldsymbol{y}}_{i}|(1+|\tilde{\boldsymbol{y}}_{p_{3}}^{d}|)}{|\tilde{\boldsymbol{y}}_{p_{3}}-\tilde{\boldsymbol{y}}_{p_{3}}^{d}|(1+|\tilde{\boldsymbol{y}}_{p_{5}}^{d}|)}>1$
@@
Appendix B Datasets and Metrics
@@
To rigorously evaluate the efficacy of our proposed loss function, we conducted comprehensive experiments across three distinct data modalities: visual data, textual data, and specialized medical corpus data (MIMIC datasets). The MIMIC datasets are particularly noteworthy for their exceptionally large label space and pronounced long-tailed distributions [12]. This long-tailed characteristic, which is especially prevalent in multi-label classification scenarios, facilitates a robust assessment of the performance of our loss function across heterogeneous data distributions. Comprehensive statistical analyses of all experimental datasets are presented in Table 3.
@@
  • MS-COCO (Microsoft Common Objects in Context) [13] consists of over 330,000 images annotated across 80 object categories, providing rich semantic information for object detection, segmentation, and captioning tasks that has significantly advanced computer vision research since its introduction by Microsoft.
@@
  • PASCAL VOC [14] contains 9,963 natural images with standardized annotations spanning 20 object categories, enabling rigorous evaluation of classification [15], detection, and segmentation algorithms in computer vision.
@@
  • NUS-WIDE [16] is a large-scale web image collection comprising approximately 269,000 Flickr images annotated with 81 concept categories and user tags, widely used as a benchmark for multi-label image classification.
@@
  • AAPD (Arxiv Academic Paper Dataset) [17] is a text corpus containing 55,840 scientific paper abstracts from arXiv with multi-label annotations across various subject categories, designed specifically for benchmarking multi-label text classification and document categorization algorithms.
@@
  • MIMIC-III  [^5] We are granted access to MIMIC-III Clinical Database (v1.4) [18] includes records labeled with expert-annotated ICD-9 codes, which identify diagnoses and procedures. We adhere to the same splits as in previous works [19], employing two settings: MIMIC-III-Full, which includes all ICD-9 codes, and MIMIC-III-50, which includes only the 50 most frequent codes.
@@
  • MIMIC-IV  [^6] We are granted access to MIMIC-IV (v2.2) [20] contains records annotated with both ICD-9 and ICD-10 codes, where each code is subdivided into sub-codes that often capture specific circumstantial details. we follow prior studies [21] and utilize four settings: MIMIC-IV-ICD9-Full, MIMIC-IV-ICD9-50, MIMIC-IV-ICD10-Full, and MIMIC-IV-ICD10-50.
Metrics. Consistent with prior research [19, 21], we report macro/micro-AUC, macro/micro-F1, and precision at K (P@$\mathcal{K}$) metrics on MIMIC datasets, where $\mathcal{K}=\{5,8\}$ for different settings. Moreover, micro/macro-F1 and mAP are used for image datasets following [6, 4, 10].
@@
Appendix C Baseline Loss Functions and Encoders
@@
This study evaluates the proposed Similarity-Dissimilarity Loss in comparison with three established baseline loss functions: (i) ALL, (ii) ANY, and (iii) MulSupCon [4], all implemented within the MSCL framework.
@@
For experimental evaluation, we employ modality-specific encoder architectures tailored to each data type. For image data, ResNet-50 [22] serves as the encoder architecture, consistent with established methodologies [6, 5, 4]. For textual data, we utilize pre-trained large language models (LLMs), specifically RoBERTa-base [23] and Llama-3.1-8B [24] with Low-Rank Adaptation (LoRA) [25]. Additionally, for the specialized task of ICD coding on MIMIC datasets, we implement PLM-ICD [26], a model specifically designed for ICD coding using LLMs.
Appendix D Implementation Details
@@
Within the MSCL framework, we implement a two-phase training method as established by Khosla [1]: (i) encoder training, wherein the model learns to generate vector representations that maximize similarity between instances of the same class while distinguishing them from other classes; and (ii) classifier training, which utilizes the trained encoder and freeze it to train the classifier. All the experiments are conducted on 2 $\times$ Nvidia A6000 48GB.
@@
In the representation training, we use a standard cosine learning rate scheduler with a $0.05$ warm-up period and set the temperature $\tau=0.07$. The projection head comprises two MLP layers with ReLU activation function and employs contrastive loss function for the training, where the projected representation $\boldsymbol{z}_{k}=Proj(Enc(\tilde{\boldsymbol{x}}_{k}))\in\mathbb{R}^{D_{P}}$. Here $h=Enc(\tilde{\boldsymbol{x}}_{k})$ denotes the encoded feature vectors and the projection dimension $D_{P}=256$. For subsequent classifier training, the projection head is removed, a linear layer is appended to the frozen encoder, and binary cross-entropy (BCE) loss is utilized for optimization.
@@
For image data, we employ ResNet-50 using stochastic gradient descent (SGD) with momentum. The input images are set up at a resolution of $224\times 224$ pixels. For text data, RoBERTa-base and Llama-3.1-8B serve as backbone encoders implemented via Huggingface platform [27]. RoBERTa configures with a dropout rate of $0.1$ and AdamW optimizer with a weight decay of $0.01$, exempting bias and LayerNorm from weight decay. Compared with full-parameter fine-tuning, we employ LoRA [25] to efficiently fine-tune large model Llama. LoRA configures with the low-rank dimension $r=16$, scaling factor $\alpha=32$ and dropout as $0.1$. There is no KV cache to save memory during training. To enhance computational efficiency, BFloat16 precision is used for the training. The hyperparameters and detailed configuration are shown our code. 777 https://github.com/guangminghuang/similarity-dissimilarity-loss
Appendix E More Experiment Results: Evaluation on Medical Domain
Table 4 reports that our method achieves superior results on MIMIC-IV-ICD9-Full compared to MIMIC-III-Full, despite both datasets employing identical ICD-9 coding standards. This marked performance differential can be attributed primarily to the more extensive training corpus available in MIMIC-IV-ICD9-Full (see in Table 3). While MIMIC-IV-ICD10-Full similarly comprises a substantial volume of clinical data, its considerably expanded label taxonomy introduces increased representational sparsity and presents additional computational and methodological challenges [21]. Moreover, the MIMIC-IV-ICD10-50 dataset demonstrates consistent performance metrics in this restricted setting, providing empirical evidence that label space dimensionality constitutes a critical determinant of model training efficacy.
Comparative analysis of model performance reveals that Llama significantly outperforms RoBERTa across evaluation metrics, a finding attributable to scaling laws of LLMs and the extensive knowledge and training corpus during the pre-training phase [28, 29, 30]. Although LLMs demonstrate considerable efficacy in domain-specific applications [31], our results indicate that PLM-ICD consistently surpasses both RoBERTa and Llama across all experimental configurations. This hierarchical performance pattern aligns with theoretical expectations, as PLM-ICD incorporates architecture and training paradigms specifically optimized for automated ICD coding tasks [26]. Despite the increasing generalization capabilities of foundation models in diverse applications, significant questions persist regarding their capacity to achieve state-of-the-art performance on highly specialized tasks, particularly within the medical domain, without substantial domain-specific training or parameter-efficient adaptation techniques [32]. Contemporary research on foundation model applications in biomedical domain has predominantly relied on specialized adaptation methods tailored to specific domain requirements. The comparative advantages of domain-specific pre-training becomes particularly evident following the development of initial foundation model architectures, as exemplified by widely implemented medical models such as Med-PaLM [33] and Med-Gemini[32].

### v5

@@ B.4 Proof of Theorem 4
Let us establish the strict inequality $\mathcal{K}_{4}^{s}\mathcal{K}_{4}^{d}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ through direct comparison. From definitions 14 and 15, we have:
$\mathcal{K}_{4}^{s}\mathcal{K}_{4}^{d}=\frac{|\tilde{\bm{y}}_{p_{4}}|}{|\tilde{\bm{y}}_{i}|}>\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}=\frac{|\tilde{\bm{y}}_{p_{3}}-\tilde{\bm{y}}_{p_{3}}^{d}|}{|\tilde{\bm{y}}_{i}|}\cdot\frac{1}{1+|\tilde{\bm{y}}_{p_{3}}^{d}|}$
@@
$\frac{|\tilde{\bm{y}}_{p_{4}}|(1+|\tilde{\bm{y}}_{p_{3}}^{d}|)}{|\tilde{\bm{y}}_{i}|(1+|\tilde{\bm{y}}_{p_{3}}^{d}|)}>\frac{|\tilde{\bm{y}}_{p_{4}}-\tilde{\bm{y}}_{p_{3}}^{d}|}{|\tilde{\bm{y}}_{i}|(1+|\tilde{\bm{y}}_{p_{3}}^{d}|)}$
@@
$|\tilde{\bm{y}}_{p_{4}}|(1+|\tilde{\bm{y}}_{p_{3}}^{d}|)>|\tilde{\bm{y}}_{p_{3}}-\tilde{\bm{y}}_{p_{3}}^{d}|$
By the cardinality constraint 22 in the theorem:
$|\tilde{\bm{y}}_{p_{3}}|(1+|\tilde{\bm{y}}_{p_{3}}^{d}|)>|\tilde{\bm{y}}_{p_{3}}-\tilde{\bm{y}}_{p_{3}}^{d}|$
@@ B.5 Proof of Theorem 5
From definitions 14 and 15, we have:
$\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}$ $=\frac{|\tilde{\bm{y}}_{p_{3}}-\tilde{\bm{y}}_{p_{3}}^{d}|}{|\tilde{\bm{y}}_{i}|}\cdot\frac{1}{1+|\tilde{\bm{y}}_{p_{3}}^{d}|}$
$\mathcal{K}_{5}^{s}\mathcal{K}_{5}^{d}$ $=\frac{1}{1+|\tilde{\bm{y}}_{p_{5}}^{d}|}$
@@
$\frac{\mathcal{K}_{5}^{s}\mathcal{K}_{5}^{d}}{\mathcal{K}_{3}^{s}\mathcal{K}_{3}^{d}}=\frac{|\tilde{\bm{y}}_{i}|(1+|\tilde{\bm{y}}_{p_{3}}^{d}|)}{|\tilde{\bm{y}}_{p_{3}}-\tilde{\bm{y}}_{p_{3}}^{d}|(1+|\tilde{\bm{y}}_{p_{5}}^{d}|)}$
@@
$|\tilde{\bm{y}}_{p_{3}}-\tilde{\bm{y}}_{p_{3}}^{d}|\leq|\tilde{\bm{y}}_{i}|$
Given the constraint 24, $|\tilde{\bm{y}}_{p_{5}}^{d}|\leq|\tilde{\bm{y}}_{p_{3}}^{d}|$, we have:
$\frac{|\tilde{\bm{y}}_{i}|(1+|\tilde{\bm{y}}_{p_{3}}^{d}|)}{|\tilde{\bm{y}}_{p_{3}}-\tilde{\bm{y}}_{p_{3}}^{d}|(1+|\tilde{\bm{y}}_{p_{5}}^{d}|)}>1$
@@
Appendix C Computing Cost Analysis
The key computation in all supervised contrastive loss functions, including ours, is the similarity between representations and the size of the positive set $\mathcal{P}(i)$ for each anchor $i$. All methods share the same core structure:
$\mathcal{L}_{i}=\frac{-1}{|\mathcal{P}(i)|}\sum_{p\in\mathcal{P}(i)}\log\frac{\exp(\bm{z}_{i}\cdot\bm{z}_{p}/\tau)}{\sum_{a\in\mathcal{A}(i)}\exp(\bm{z}_{i}\cdot\bm{z}_{a}/\tau)}$
Our method introduces a reweighting factor $\mathcal{K}_{i,p}^{s}\mathcal{K}_{i,p}^{d}\in[0,1]$ only in the numerator of the positive terms:
$\mathcal{L}_{i}^{\text{our}}=\frac{-1}{|\mathcal{P}(i)|}\sum_{p\in\mathcal{P}(i)}\log\frac{\mathcal{K}_{i,p}^{s}\mathcal{K}_{i,p}^{d}\exp(\bm{z}_{i}\cdot\bm{z}_{p}/\tau)}{\sum_{a\in\mathcal{A}(i)}\exp(\bm{z}_{i}\cdot\bm{z}_{a}/\tau)}$
Importantly:
  • The denominator (negative set) is unchanged across ALL, ANY, MulSupCon, and our method.
  • The positive set $\mathcal{P}(i)$ for our method is structurally identical to ANY, typically larger than ALL but smaller or equal to MulSupCon, which forms multiple sets.
  • The computation of $\mathcal{K}_{i,p}^{s}=\frac{|\mathcal{S}\cap\mathcal{T}|}{|\mathcal{S}|}$ and $\mathcal{K}_{i,p}^{d}=\frac{1}{1+|\mathcal{T}\setminus(\mathcal{S}\cap\mathcal{T})|}$ involves simple set operations on one-hot label vectors, which are efficient and linear in the number of labels.
  • No additional encoder forward passes or model parameters are introduced.
Empirically, memory and runtime cost for the Similarity-Dissimilarity Loss:
  • Does not increase memory usage, as it does not alter the encoder architecture or increase feature dimensionality.
  • Adds minimal runtime cost, primarily due to lightweight scalar computations per positive pair (i.e., computing $\mathcal{K}_{i,p}^{s}\mathcal{K}_{i,p}^{d}$).
In contrast, methods like MulSupCon may require repeated label-wise positive sets and multiple forward loss computations per label, depending on implementation. Thus, our method maintains comparable or better efficiency while offering a more principled weighting mechanism.
Appendix D Datasets and Metrics
@@
To rigorously evaluate the efficacy of our proposed loss function, we conducted comprehensive experiments across three distinct data modalities: visual data, textual data, and specialized medical corpus data (MIMIC datasets). The MIMIC datasets are particularly noteworthy for their exceptionally large label space and pronounced long-tailed distributions (Huang et al., 2024a). This long-tailed characteristic, which is especially prevalent in multi-label classification scenarios, facilitates a robust assessment of the performance of our loss function across heterogeneous data distributions. Comprehensive statistical analyses of all experimental datasets are presented in Table 2.
@@
  • MS-COCO (Microsoft Common Objects in Context) (Lin et al., 2014) consists of over 330,000 images annotated across 80 object categories, providing rich semantic information for object detection, segmentation, and captioning tasks that has significantly advanced computer vision research since its introduction by Microsoft.
@@
  • PASCAL VOC (Everingham et al., 2010) contains 9,963 natural images with standardized annotations spanning 20 object categories, enabling rigorous evaluation of classification (Huang et al., 2019), detection, and segmentation algorithms in computer vision.
@@
  • NUS-WIDE (Chua et al., 2009) is a large-scale web image collection comprising approximately 269,000 Flickr images annotated with 81 concept categories and user tags, widely used as a benchmark for multi-label image classification.
@@
  • AAPD (Arxiv Academic Paper Dataset) (Yang et al., 2018) is a text corpus containing 55,840 scientific paper abstracts from arXiv with multi-label annotations across various subject categories, designed specifically for benchmarking multi-label text classification and document categorization algorithms.
@@
  • MIMIC-III  [^6] We are granted access to MIMIC-III Clinical Database (v1.4) (Johnson et al., 2016) includes records labeled with expert-annotated ICD-9 codes, which identify diagnoses and procedures. We adhere to the same splits as in previous works (Mullenbach et al., 2018), employing two settings: MIMIC-III-Full, which includes all ICD-9 codes, and MIMIC-III-50, which includes only the 50 most frequent codes.
@@
  • MIMIC-IV  [^7] We are granted access to MIMIC-IV (v2.2) (Johnson et al., 2020) contains records annotated with both ICD-9 and ICD-10 codes, where each code is subdivided into sub-codes that often capture specific circumstantial details. we follow prior studies (Nguyen et al., 2023) and utilize four settings: MIMIC-IV-ICD9-Full, MIMIC-IV-ICD9-50, MIMIC-IV-ICD10-Full, and MIMIC-IV-ICD10-50.
[…]
@@
Appendix E Baseline Loss Functions and Encoders
@@
This study evaluates the proposed Similarity-Dissimilarity Loss in comparison with three established baseline loss functions: (i) ALL, (ii) ANY, and (iii) MulSupCon (Zhang & Wu, 2024), all implemented within the MSCL framework.
@@
For experimental evaluation, we employ modality-specific encoder architectures tailored to each data type. For image data, ResNet-50 (He et al., 2016) serves as the encoder architecture, consistent with established previous studies (He et al., 2020; Chen et al., 2020; Zhang & Wu, 2024), which is a main and common encoder in previous works. For textual data, we utilize pre-trained large language models (LLMs), specifically RoBERTa-base (Liu et al., 2019) and Llama-3.1-8B (Grattafiori et al., 2024) with Low-Rank Adaptation (LoRA) (Hu et al., 2022). Additionally, for the specialized task of ICD coding on MIMIC datasets, we implement PLM-ICD (Huang et al., 2022), a model specifically designed for ICD coding using LLMs.
Appendix F Implementation Details
@@
Within the MSCL framework, we implement a two-phase training method as established by Khosla (Khosla et al., 2020): (i) encoder training, wherein the model learns to generate vector representations that maximize similarity between instances of the same class while distinguishing them from other classes; and (ii) classifier training, which utilizes the trained encoder and freeze it to train the classifier. All the experiments are conducted on 2 $\times$ Nvidia A6000 48GB.
@@
In the representation training, we use a standard cosine learning rate scheduler with a $0.05$ warm-up period and set the temperature $\tau=0.07$. The projection head comprises two MLP layers with ReLU activation function and employs contrastive loss function for the training, where the projected representation $\bm{z}_{k}=Proj(Enc(\tilde{\bm{x}}_{k}))\in\mathbb{R}^{D_{P}}$. Here $h=Enc(\tilde{\bm{x}}_{k})$ denotes the encoded feature vectors and the projection dimension $D_{P}=256$. For subsequent classifier training, the projection head is removed, a linear layer is appended to the frozen encoder, and binary cross-entropy (BCE) loss is utilized for optimization.
@@
For image data, we employ ResNet-50 using stochastic gradient descent (SGD) with momentum. The input images are set up at a resolution of $224\times 224$ pixels. For text data, RoBERTa-base and Llama-3.1-8B serve as backbone encoders implemented via Huggingface platform (Wolf et al., 2020). RoBERTa configures with a dropout rate of $0.1$ and AdamW optimizer with a weight decay of $0.01$, exempting bias and LayerNorm from weight decay. Compared with full-parameter fine-tuning, we employ LoRA (Hu et al., 2022) to efficiently fine-tune large model Llama. LoRA configures with the low-rank dimension $r=16$, scaling factor $\alpha=32$ and dropout as $0.1$. There is no KV cache to save memory during training. To enhance computational efficiency, BFloat16 precision is used for the training. The hyperparameters and detailed configuration are shown our code.  [^8] https://github.com/anonymous/similarity-dissimilarity-loss
Appendix G More Experiment Results and Analysis
G.1 Detailed Results and Analysis on Image
Figure 2(a) illustrates the comparison between Similarity-Dissimilarity Loss and MulSupCon as measured by micro- and macro-F1 metrics. The results indicate that our method yields substantially greater improvements in macro-F1 compared to micro-F1 across all image datasets. Specifically, macro-F1 increases by $5.7\%$ on MS-COCO and $5.8\%$ on NUS-WIDE, whereas micro-F1 exhibits more modest improvements of $2.9\%$ and $2.0\%$, respectively. Macro-F1 assigns equal importance to each class regardless of its frequency, rendering it particularly appropriate for evaluating performance on imbalanced datasets where minority class prediction accuracy is critical (LeCun et al., 2015; Zhang et al., 2023). In contrast, micro-F1 places more considerable weight on classes with more samples, making it more appropriate when larger classes should have a more potent influence on the overall score (LeCun et al., 2015; Lin et al., 2017). Multi-label classification inherently faces more pronounced challenges with long-tailed distributions than single-label classification due to exponential output space complexity, intricate label co-occurrence patterns, and high annotation costs (Zhang et al., 2023). The observed superior improvement in macro-F1 metrics provides compelling evidence that our method demonstrates exceptional efficacy in addressing long-tailed distribution challenges, a capability particularly crucial in multi-label scenarios.
However, on the PASCAL dataset, our method demonstrates mere marginal improvements, with gains of $0.88/0.84/0.17$ in micro/macro-F1/mAP, respectively. This limited enhancement can be attributed to the structural characteristics of PASCAL, wherein the average number of labels per instance is approximately $1.5$ (as detailed in Table 2), causing the task to approximate single-label classification, particularly when the batch size is limited (Khosla et al., 2020). Consequently, loss functions specifically designed for multi-label scenarios exert minimal influence on model performance under these conditions. As Audibert et al. (2024) have demonstrated, the cardinality of the label space constitutes a significant determinant of model efficacy within MSCL .
Furthermore, Figure 2(b) reveals that the standard deviation across four methods for PASCAL equals $0.58/0.64/0.27$ in micro/macro-F1/mAP, which are considerably lower than the corresponding standard deviations observed for the MS-COCO and NUS-WIDE. This statistical finding suggests that the efficacy of specialized multi-label loss functions diminishes significantly when the average label cardinality per instance approaches $1$ in MSCL. This finding further corroborates our theoretical analysis and hypothesis in the Section 2, wherein Similarity-Dissimilarity Loss degenerates to single-label scenarios (see Equation 17).
G.2 Evaluation on Text
We further evaluate our method on general text data, and the results demonstrate that our proposed loss function consistently surpasses baseline methods for both RoBERTa and Llama models across all metrics on the AAPD dataset (See Table 3). In contrast to the significant performance gains observed on image data, Similarity-Dissimilarity Loss achieves more modest enhancements of $0.90/1.79$ in micro/macro-F1 scores on RoBERTa, and $0.89/1.84$ on Llama. This attenuated performance differential can be attributed to the extensive knowledge already encoded within LLMs through their comprehensive pre-training paradigms (Yang et al., 2024).
Method RoBERTa Llama
Micro-F1 Macro-F1 Micro-F1 Macro-F1
ALL 73.23 59.41 74.32 60.47
ANY 72.31 58.55 73.41 59.63
MulSupCon 73.64 60.52 74.72 61.58
Ours 74.54 62.31 75.61 63.42
G.3 Evaluation on Medical Domain